OpenStudy (anonymous):

Evaluate without using a calculator: cos(sin^-1(2/7))

OpenStudy (anonymous):

\[\cos(\sin ^{-1}(\frac{ 2 }{ 7 }))\] heres a cleaner version of the equation

OpenStudy (anonymous):

my guess would to make a triangle. opp/hyp where opp=2 and hyp=7

OpenStudy (anonymous):

and then?

jimthompson5910 (jim_thompson5910):

theta = arcsin(2/7) sin(theta) = 2/7 so,